News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 3 September 2021 9.00am EDT

During the last week the observing schedule was interrupted on Sept 1 at 6:34 a.m. due to a failure indication in the LETG software logic. Preliminary assessment is that the grating inserted correctly, but 0.3% faster than the software expected. This resulted in contact with the mechanical stop, which left the insert microswitch in the wrong position, leading the OBC software to declare the move a failure. This prevented further OTG motions and resulted in several misconfigured observations before SCS 107 could be executed to stop the science loads. Observations impacted by the interruption will be rescheduled in a future week.

A real-time procedure was performed on Sep 1 which cleared the "last move failed flag" which is required to move the gratings. A second real-time procedure was completed on Sep 2 to retract the LETG in preparation for resuming science. Science loads resumed Sep 2 at 4:37 p.m EDT.

A press release was issued Aug 31 describing observations of the black hole system MG B2016+112. Using Chandra observations and radio based modeling, observers have concluded that the three X-ray sources they detected from the MG B2016+112 system must have resulted from the lensing of two distinct objects. These two X-ray-emitting objects are likely a pair of growing supermassive black holes or a growing supermassive black hole and its jet. For details see:

The schedule of targets for the next week is shown below and includes observations of Jupiter coordinated with HST.

        Radiation Belts                     Sep  5
        SDSSJ231838.49-1030    ACIS-S
        [DAG2016]PSZ1-ARCG     ACIS-S       Sep  6
        SDSSJ231838.49-1030    ACIS-S
        Radiation Belts                     Sep  7
        NGC3998                ACIS-S
        SDSSJ225438.30+2327    ACIS-S
        NGC3998                ACIS-S
        3CR_455                ACIS-S       Sep  8
        Jupiter                HRC-S
        Jupiter                HRC-S
        ArLac (21)             HRC-I
        SNR0519-69.0           ACIS-S       Sep  9
        Radiation Belts
        SNR0519-69.0           ACIS-S
        HD4915                 HRC-S        Sep 10
        A1351                  ACIS-I
        UGC12237               ACIS-S
        PKS2300-18             ACIS-S       Sep 11
        3FGLJ1016.5-6034       ACIS-I
        HD4915                 HRC-S
        NGC3998                ACIS-S
        Radiation Belts                     Sep 12
        NGC3998                ACIS-S


Use of the gratings is suspended pending a detailed review of the anomaly.
All other spacecraft subsystems continued to support nominal operations.

Return to Status Reports